# Hyperband-based Bayesian Optimization for Black-box Prompt Selection

### v1

@@ 3.2 Structural-aware Deep Kernel
To learn a lower-dimensional representation of the embedded prompts aligned to the downstream task, we propose to use a deep kernel (Wilson et al., 2016). We design a feature extractor, $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{p}$, $p<<d$ to learn a flexible kernel transformation function: $k(\phi(\mathbf{z},\mathbf{w}),\phi(\mathbf{z}^{\prime},\mathbf{w})|\mathbf{\theta})$, where $\mathbf{\theta}$ and $\mathbf{w}$ are the parameters of the kernel and the extractor.
Given that prompts are composed of two distinct building blocks, we hypothesize that embedding these components separately would enhance the deep kernel GP (DK-GP)’s ability to exploit both structural and semantic differences between them. For instance, instructions tend to be relatively similar across prompts, while few-shot exemplars can vary significantly due to their diverse input-output pairs and varying orders.
To address this, we propose learning a structure-aware latent representation of prompts. This approach involves embedding the instructions $i\in\mathcal{I}$ and few-shot exemplars $e\in\mathcal{E}$ of prompts separately. We proceed to use two distinct feed forward neural networks consisting of two layers with ReLU activation functions for the embedded instructions and exemplars: $\phi_{enc(\cdot)}$: Lin(d, 64) $\rightarrow$ ReLU() $\rightarrow$ Lin(64, 32) $\rightarrow$ ReLU(). After having processed the instructions and exemplars separately, we concatenate the outputs and proceed with another feed forward neural network to learn a joint latent representation: $\phi_{\left(\phi_{enc(i)},\phi_{enc(e)}\right)}$: Lin(32 $\cdot$ 2, 32) $\rightarrow$ ReLU() $\rightarrow$ Lin(32, 10).
During training the GP we obtain both the optimal kernel parameter values, as well as the parameters of the neural network feature extractors by optimizing the log marginal likelihood criterion:
$\hat{\mathbf{\theta}},\hat{\mathbf{w}}=\operatorname*{arg\,max}_{\mathbf{\theta},\mathbf{w}}-\mathbf{v}^{\intercal}\mathbf{K}(\mathbf{\theta},\mathbf{w})^{-1}\mathbf{v}-\log|\mathbf{K}(\mathbf{\theta},\mathbf{w})|$ (6)
In Section 3.4 we illustrate how to make use of our structural-aware DK-GP during optimization. First, we describe how we achieve query-efficiency.
@@ 3.3 Hyperband for Multi-Fidelity Scheduling
To boost query-efficiency identified as one of the limitation of the existing works (EASE and MIPROv2), we want to terminate the evaluation of poor performing prompts early saving time (and cost) during the evaluation process. Similarly to Shi et al. (2024) we model the number of validation instances prompts are evaluated on as a fidelity parameter. Full-fidelity methods evaluate prompts on all validation instances, while multi-fidelity methods adaptively schedule evaluations on varying numbers of instances. TRIPLE implements Successive Halving (SH; Karnin et al. 2013) to schedule the number of validation instances. In contrast, we use the Hyperband (HB; Li et al. 2018) algorithm because it will not evaluate all prompts and hedges against a poorly configured SH as we explain below.
Given a total budget of $B$ LLM calls to evaluate prompts, SH allocates a budget of $b=B/(|\mathcal{P}|\log_{2}{(|\mathcal{P}|)}$ to each prompt (see further details in Appendix C). After having evaluated the prompts on $b$ validation instances, the lower half of bad performing prompts is discarded and the process repeats, doubling the number of calls for the remaining prompts in the next stage, until a single prompt remains.
This strategy is affected by the ”budget vs. number of configurations” dilemma (Li et al., 2018), since, at the beginning of the algorithm it is not clear if it should evaluate many (by default all) prompts on few instances (good exploration but noisy performance estimates) or few prompts on many instances (less exploration but accurate performance estimates). However, when many prompts need to be evaluated within a limited total budget, SH’s initial budget will be low which risks discarding a prompt based on noisy performance estimation. In contrast, we propose to use HB since it hedges against a poor choice of the number of starting prompts and budget by repeatedly running SH in different brackets using different numbers of starting prompts and starting budgets. This allows HB to be robust under various scenarios without knowing the optimal resource allocation beforehand making it ideal for prompt selection.
To adapt HB to prompt selection we make the following design choices: (1) We extend previous evaluations when advancing stages within a bracket, ensuring validation instances of higher stages subsume those of lower stages and (2) return the prompt with the lowest validation error among those evaluated on the entire validation set. (3) In addition, we combine HB with our structural-aware DK-GP by employing a sequential proposal mechanism for candidate prompts in each bracket, which we outline next.
@@
0:  $n_{\mathrm{valid}}$, $b_{\min}$: lower limit to #validation instances, $\eta$: halving parameter
@@ 3.4 HbBoPs
HbBoPs combines HB with our structural-aware DK-GP. While the vanilla HB algorithm for prompt selection samples prompts uniformly at random, HbBoPs replaces the random proposal mechanism of HB with a sequential BO proposal (highlighted in gray) similar to the approach proposed by Falkner et al. (2018) for Hyperparameter Optimization. During the execution of HB, HbBoPs trains the GP on a subset of the design data ${\mathcal{D}_{t}}_{|b}$ for a given fidelity-level $b$. We use the highest fidelity $b$ for which ”enough” observations have been collected. This design choice stems from the observation that validation errors are estimated more accurately with more instances (see Appendices C and B).
After having trained the GP on ${\mathcal{D}_{t}}_{|b}$, we obtain the next candidate prompt ($p_{t+1}$) for evaluation based on the Expected Improvement (Jones et al., 1998) as acquisition function:
@@
where, given the incumbent (the best performing prompt on the highest fidelity evaluated yet) and its validation error $v_{\min,b}$ at the given fidelity level $b$, the Expected Improvement tells us how much we can expect the given candidate prompt to improve over this incumbent given the posterior predictive distribution as described in Equation (5) of the GP. The full algorithm is described in Algorithm 1.
@@
Method Fidelity Level Efficiency Surrogate Model Bandit Algorithm Prompt Representation sample query RS Full - - - - $p$ Vanilla BO Full - vGP - $enc(p)$ HDBO Full - GP (Hvarfner et al., 2024) - $enc(p)$ BOPCA Full - vGP - $\text{PCA}(enc(p))$ (Zhang et al., 2024) EASE (Wu et al., 2024) Full - NN NUCB $enc(p)$ MIPROv2 (Opsahl-Ong et al., 2024) Full - TPE - $\text{ID}_{i}\,\text{ID}_{e}$ TRIPLE-SH (Shi et al., 2024) Multi - - SH $p$ TRIPLE-GSE (Shi et al., 2024) Multi - LM/GLM GSE $enc(p)$ HbBoPs (ours) Multi DK-GP HB $enc(i),enc(e)$
@@ 4.1 Benchmark Tasks
To benchmark HbBoPs we select ten tasks commonly used for LLM evaluation (Zhou et al., 2022; Lin et al., 2023; Chen et al., 2023; Wu et al., 2024; Shi et al., 2024). AI2’s Reasoning Challenge (ARC) (Clark et al., 2018): multiple choice question answering problems; Grade School Math 8K (Cobbe et al., 2021): math problems that take between two and eight steps to solve; 8 Tasks from the BBII subset of the BIG-bench and instruction induction benchmarks (Srivastava et al., 2022; Honovich et al., 2022) used in Zhou et al. (2022); Wu et al. (2024); Shi et al. (2024): antonyms, larger animal, negation, second word letter, sentiment, object counting, orthography starts with, and word unscrambling. Task statistics are reported in Table 5 in Appendix D.
@@ 4.2 Methods
We compare HbBoPs against full-fidelity and multi-fidelity methods described in Table 1. Additional details of the methods are reported in Section 6.2 and Appendix D. All methods that rely on embeddings of prompts use BERT’s [CLS] token embedding. All full-fidelity BO methods (vanilla BO, HDBO, BOPCA) use an ARD Matérn $\frac{5}{2}$ kernel and Expected Improvement as acquisition function and normalize inputs to the unit cube and standardize outputs. HDBO is a simply but well performing high-dimensional BO algorithm and uses adjusted priors on kernel and likelihood parameters as described in Hvarfner et al. (2024). BOPCA uses a ten component PCA inspired by Zhang et al. (2024). We run MIPROv2, EASE’s NUCB ($\nu=0.1$), TRIPLE-SH and TRIPLE-GSE as implemented in their publicly available code bases. All full-fidelity methods use the same initial design of ten prompts sampled uniformly at random. HbBoPs uses an ARD Matérn $\frac{5}{2}$ kernel, normalizes inputs and standardizes outputs. We train the DK-GP on the highest fidelity level for which at least 4 observations are available. To optimize the log marginal likelihood described in Equation (6), we use AdamW (Loshchilov, 2017) with learning rate $=0.01$, maximum number of epochs $=3,000$ epochs, and early termination with a patience $=10$. Within the HB schedule we use a lower limit to the number of validation instances $b_{\min}=10$ and a halving parameter $\eta=2.0$.
@@ 4.3 Experimental Protocol
For each task, we generate a search space $\mathcal{P}$ of candidate prompts by combining five task-specific instructions with 50 few-shot exemplars. Instructions are created using APE’s forward mode (Zhou et al., 2022), where Claude 3 Sonnet (Anthropic, 2024) generates instructions based on ten input-output samples from each task’s training set. For exemplars, we sample 25 sets of five input-output instances from the training set of each task, then permute each set twice to create 50 exemplar tuples, allowing assessment of example ordering effects. The final 250 prompts are formed by the Cartesian product of instructions and exemplars.
For LLMs, we use Claude 3 Haiku (Anthropic, 2024), LLAMA3 8B Instruct (Dubey et al., 2024), and Mistral 7B Instruct (Jiang et al., 2023).
For each benchmark scenario (a given task and LLM), we run each method for a total budget of 25 full-fidelity evaluations (i.e., being allowed as many LLM calls as 25 prompts evaluated on all validation instances would require for a given task) to mimic a budget constrained scenario. We repeat each method run 30 times. We evaluate prompts using the loss function described in Equation (1) that is based on the exact match scoring function.
@@
Figure 1: Anytime normalized error of the best prompt found by each method, averaged over different benchmark scenarios. Lower is better. Log scale. Ribbons represent standard errors.
@@ 5 Results
In this section, we discuss the experimental results. We report the validation and test errors computed on the best prompt identified by each method given a specific budget. For instance, given GSM8k with a validation set of 1319 instances, a budget of $0.25$ means that we report the results of the methods after performing $int(0.25\cdot 25\cdot 1319)=8244$ LLM calls. Therefore, full-fidelity methods always start after having executed a fraction of $\nicefrac{{1}}{{25}}$ total LLM calls.
@@ 5.1 Analysis of overall performance
We start by analyzing the overall performance of the methods averaged across all the benchmark tasks and LLMs. To allow for averaging results over different benchmarks and models, we normalize validation and test errors for each benchmark scenario by the performance of the worst and best prompt. Figure 1 visualizes the results.
Analysis of the figures reveals that HbBoPs outperforms all full-fidelity and multi-fidelity methods, particularly in terms of anytime performance on both the validation and test sets.
Beginning with an analysis of test performance at full budget (i.e., a fraction of LLM calls equal to 1.0), we can see that our HbBoPs on average outperforms all full-fidelity and multi-fidelity approaches with an average normalized test error of 0.150. In detail, we observe that all full-fidelity methods surpass the RS baseline (0.214) with the following errors: Vanilla BO (0.211), MIPROv2 (0.198), EASE (0.195), BOPCA (0.192), and HDBO (0.185). However, they all have higher error values than HbBoPs (0.1500). Additionally, HbBoPs also outperforms all the multi-fidelity methods. Although both TRIPLE-GSE (0.158) and TRIPLE-SH (0.159) exhibit superior performance compared to their best-in-class full-fidelity counterpart, i.e., HDBO (0.185), they on average have identified prompts that yield error values higher than the ones obtained for HbBoPs’s prompts.
Looking at the anytime performance with a more limited budget, e.g., a fraction of 0.25 LLM calls, we can confirm HbBoPs’s improvements over the baselines. Indeed, HbBoPs on average outperforms HDBO, the best full-fidelity method, by approximately 35% and TRIPLE-SH, the best multi-fidelity method, by 24%.
5.2 Analysis of the performance for each LLMs
Table 2: Median relative validation and test improvement of HbBoPs over TRIPLE-SH over the ten benchmarks for each LLM for different fractions of total LLM calls.

### v2

@@ 3.2 Structural-aware Deep Kernel
To learn a lower-dimensional representation of the embedded prompts aligned with the downstream task, we propose to use a deep kernel (Wilson et al., 2016) within the GP. We design a feature extractor, $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{p}$, $p\ll d$ to learn a flexible kernel transformation function $k(\phi(\mathbf{z},\mathbf{w}),\phi(\mathbf{z}^{\prime},\mathbf{w})|\mathbf{\theta})$, where $\mathbf{\theta}$ and $\mathbf{w}$ are the parameters of the kernel and respectively the extractor.
Given that prompts consist of two distinct components (instructions and few-shot exemplars), we hypothesize that embedding these components separately can improve the deep kernel GP’s (DK-GP) ability to use both structural and semantic differences effectively. Instructions typically exhibit consistent patterns across prompts, whereas few-shot exemplars are more variable due to diverse input-output pairs and flexible ordering. To address this, we propose learning a structural-aware latent representation of prompts. Our approach involves embedding the instructions $i\in\mathcal{I}$ and few-shot exemplars $e\in\mathcal{E}$ separately. Each component embedding is processed through distinct feed-forward neural networks, each consisting of two fully connected layers with ReLU activations, as defined below:
ϕenc()::subscriptitalic-$\phi_{\text{enc}(\cdot)}:$
Lin(d, 64) $\rightarrow$ ReLU() $\rightarrow$ Lin(64, 32) $\rightarrow$ ReLU()
After processing the instructions and exemplars independently, we concatenate their outputs and input the combined representation into another feed-forward neural network to learn a joint latent representation:
$\phi_{\left(\phi_{\text{enc}(i)},\phi_{\text{enc}(e)}\right)}:$
[…]
During training the GP, we obtain both the optimal kernel parameters and the parameters of the neural network feature extractor $\phi$ by optimizing the log marginal likelihood criterion (up to constants):
$\hat{\mathbf{\theta}},\hat{\mathbf{w}}=\operatorname*{arg\,max}_{\mathbf{\theta},\mathbf{w}}-\mathbf{v}^{\intercal}\mathbf{K}_{t}(\mathbf{\theta},\mathbf{w})^{-1}\mathbf{v}-\log|\mathbf{K}_{t}(\mathbf{\theta},\mathbf{w})|$ (6)
An additional advantage of using a deep kernel GP is its ability to model non-stationary functions, as the feature extractor learns input-dependent transformations that allow the surrogate to capture varying levels of smoothness across the input space, which standard stationary kernels in vanilla GPs cannot accomplish (Li et al., 2024). Before illustrating how we use our structural-aware DK-GP during optimization to achieve sample-efficiency via a BO proposal, we explain how we ensure query-efficiency via Hyperband.
@@ 3.3 Hyperband for Multi-Fidelity Scheduling
To improve query-efficiency identified as one of the key limitations of existing works, we want to terminate the evaluation of poor-performing prompts early, saving cost during the evaluation process. Similarly to TRIPLE (Shi et al., 2024), we model the number of validation instances prompts are evaluated on as a fidelity parameter. Full-fidelity methods evaluate prompts on all validation instances, while multi-fidelity methods adaptively schedule evaluations on varying numbers of instances. TRIPLE implements Successive Halving (SH; Karnin et al. 2013) as a multi-fidelity scheduler. In contrast, we use Hyperband (HB; Li et al. 2018) as it will generally evaluate fewer prompts and hedges against a poorly configured SH as explained below.
Given a total budget of $B$ LLM calls to evaluate prompts on validation instances, SH allocates a budget of $b=B/(|\mathcal{P}|\log_{2}{(|\mathcal{P}|)})$ to each prompt (see details in Appendix C). After having evaluated the prompts on $b$ validation instances, the lower half of bad performing prompts is discarded, and the process repeats, doubling the number of calls for the remaining prompts in the next stage, until a single prompt remains.
This strategy is affected by the “budget vs. number of configurations” dilemma (Li et al., 2018), since, at the beginning of the algorithm, it is not clear if we should evaluate many (by default all) prompts on few instances (good exploration but noisy performance estimates) or few prompts on many instances (less exploration but accurate performance estimates). When many prompts need to be evaluated with a limited total budget, SH’s initial budget is low, which risks discarding a prompt based on noisy performance estimation. HB in contrast hedges against a poor choice of the number of starting prompts and their budget by repeatedly running SH in different brackets with different numbers of starting prompts and starting budgets. This results in HB being robust under various scenarios without knowing the optimal resource allocation, making it ideal for prompt selection.
To tailor HB to prompt selection, we make the following design decisions (for an ablation study, see Appendix E): (1) We extend previous evaluations when advancing stages within a bracket, ensuring validation instances of higher stages subsume those of lower stages, and (2) return the prompt with the lowest validation error among those evaluated on the entire validation set.
Finally, we combine HB with our structural-aware DK-GP by employing a sequential BO proposal mechanism for candidate prompts in each bracket, which we outline next.
@@
0:  $n_{\mathrm{valid}}$, $b_{\min}$ (lower limit to #validation instances), $\eta$ (halving parameter)
@@ 3.4 HbBoPs
HbBoPs combines HB with our structural-aware DK-GP. While the vanilla HB algorithm for prompt selection samples prompts uniformly at random, HbBoPs replaces the random proposal mechanism of HB with a sequential BO proposal (highlighted in gray in Algorithm 1). This is similar to the approach proposed by Falkner et al. (2018) for hyperparameter optimization. During the execution of HB, HbBoPs trains the GP on a subset of the design data ${\mathcal{D}_{t}}_{|b}$ for a given fidelity-level $b$. We use the highest fidelity $b$ for which “enough” observations are available. This design decision stems from the observation that validation errors are estimated more accurately with more instances (see Appendices B and C). Importantly, we train the GP entirely online during the selection process and are not relying on a pre-trained surrogate model.
After training the GP on ${\mathcal{D}_{t}}_{|b}$, we select the next candidate prompt $p_{t+1}$ by maximizing the Expected Improvement (EI; Kushner 1964; Mockus et al. 1978; Jones et al. 1998) acquisition function:
@@
In words, given the incumbent (the best-performing prompt evaluated at the highest fidelity level so far) and its validation error $v_{\min,b}$ at fidelity level $b$, the EI quantifies the expected improvement of a candidate prompt over the incumbent (considering only actual improvement due to the $\max$ operator), based on the GP’s posterior predictive distribution (Equation (5)). Since our search space is given by a finite set of candidate prompts, we can evaluate the EI exhaustively for all candidate prompts.
@@
Method Fidelity Level Efficiency Surrogate Model Bandit Algorithm Prompt Representation sample query RS Full - - - - $p$ Vanilla BO Full - vanilla GP - $enc(p)$ HDBO Full - GP (Hvarfner et al., 2024) - $enc(p)$ BOPCA Full - vanilla GP - $\text{PCA}(enc(p))$ (Zhang et al., 2024) EASE (Wu et al., 2024) Full - NN NUCB $enc(p)$ MIPROv2 (Opsahl-Ong et al., 2024) Full - TPE - $\text{ID}_{i}\,\text{ID}_{e}$ TRIPLE-SH (Shi et al., 2024) Multi - - SH $p$ TRIPLE-GSE (Shi et al., 2024) Multi - LM/GLM GSE $enc(p)$ HbBoPs (ours) Multi structural-aware DK-GP HB $enc(i),enc(e)$
@@ 4.1 Benchmark Tasks
We benchmark HbBoPs on ten tasks commonly used for LLM evaluation (Zhou et al., 2023; Lin et al., 2024; Chen et al., 2024; Wu et al., 2024; Shi et al., 2024). AI2’s Reasoning Challenge (ARC) (Clark et al., 2018): multiple-choice Q&A problems; Grade School Math 8K (Cobbe et al., 2021): math problems taking between two and eight steps to solve; Eight Tasks from the BBII subset of the BIG-bench and instruction induction benchmarks (Srivastava et al., 2023; Honovich et al., 2023) used in Zhou et al. (2023); Wu et al. (2024); Shi et al. (2024): antonyms, larger animal, negation, second word letter, sentiment, object counting, orthography starts with, and word unscrambling. Task statistics are reported in Table 5 in Appendix D.
@@ 4.2 Methods
We compare HbBoPs against full-fidelity and multi-fidelity methods described in Table 1. Additional details on the methods are reported in Section 6 and Appendix D. RS is a simple random search. All methods that rely on embeddings of prompts use BERT’s [CLS] token embedding to ensure fair comparison. All full-fidelity BO methods (vanilla BO, HDBO, BOPCA) use an ARD Matérn $\nicefrac{{5}}{{2}}$ kernel and Expected Improvement as acquisition function and normalize inputs to the unit cube and standardize outputs to have zero mean and unit variance. HDBO is a simple but well-performing high-dimensional BO algorithm using adjusted priors on GP kernel and likelihood parameters, as described in Hvarfner et al. (2024). BOPCA uses a ten component PCA inspired by Zhang et al. (2024). We run MIPROv2, NUCB (Zhou et al., 2020) as used by EASE ($\nu=0.1$), TRIPLE-SH and TRIPLE-GSE as implemented in their publicly available code bases. All full-fidelity methods use the same initial design of ten prompts sampled uniformly at random. HbBoPs uses an ARD Matérn $\nicefrac{{5}}{{2}}$ kernel, normalizes inputs to the unit cube and standardizes outputs. We always train the DK-GP on the highest fidelity level for which at least four observations are available. To optimize the log marginal likelihood in Equation (6), we use AdamW (Loshchilov & Hutter, 2019) with learning rate $=0.01$, maximum number of epochs $=3000$, and early termination with a patience $=10$. Within the HB schedule, we use a lower limit on the number of validation instances $b_{\min}=10$ and a halving parameter $\eta=2.0$.
@@ 4.3 Experimental Protocol
For each task, we generate a search space $\mathcal{P}$ of candidate prompts by combining five task-specific instructions with 50 few-shot exemplars. Instructions are generated using APE’s forward mode (Zhou et al., 2023), where Claude 3 Sonnet (Anthropic, 2024) generates instructions based on ten input-output samples from each task’s training set. For exemplars, we sample 25 sets of five input-output instances from the training set of each task, then permute each set twice to create 50 exemplar tuples, allowing assessment of example ordering effects. We fix the five-shot setting throughout our experiments to ensure consistency with prior work (Wu et al., 2024). The final $|\mathcal{P}|=250$ prompts are formed by the Cartesian product of instructions and exemplars.
For LLMs, we use Claude 3 Haiku (Anthropic, 2024), LLAMA3 8B Instruct (Llama Team, AI @ Meta, 2024), and Mistral 7B Instruct (Jiang et al., 2023).
For each benchmark scenario (a given task and LLM), we run each method for a total budget of 25 full-fidelity evaluations (i.e., being allowed as many LLM calls as 25 prompts evaluated on all validation instances require for a given task) to mimic a budget-constrained scenario. We use the number of LLM calls as our cost metric, rather than actual monetary cost, since LLM calls are a directly interpretable and model-agnostic measure of cost that allows for aggregating over different benchmark scenarios. We repeat each method run 30 times on each benchmark scenario. We evaluate prompts using the loss function described in Equation (1) which is based on the exact match scoring function.
@@
Figure 1: Normalized error (log scale) of the best prompt per method, averaged over benchmarks. Lower is better. Ribbons represent SE.
@@ 5 Results
We report the validation and test errors computed for the best prompt identified by each method given a specific budget. For instance, given GSM8k with a validation set of 1319 instances, a budget of $0.25$ means that we report the results of the methods after performing $\lceil 0.25\cdot 25\cdot 1319\rceil=8244$ LLM calls. Therefore, full-fidelity methods always start after having executed a fraction of $\nicefrac{{1}}{{25}}$ total LLM calls.
@@ 5.1 Analysis of Overall Performance
We start by analyzing the overall performance of the methods averaged across all benchmark tasks and LLMs. To allow for averaging results, we normalize validation and test errors for each benchmark scenario by the performance of the worst and best prompt. Figure 1 visualizes the results. We observe that HbBoPs outperforms all full-fidelity and multi-fidelity methods, particularly in terms of anytime performance on both the validation and test set.
Beginning with an analysis of test error at full budget (i.e., a fraction of LLM calls equal to 1.0), we can see that our HbBoPs on average outperforms all full-fidelity and multi-fidelity approaches with an average normalized test error of 0.150. In detail, we observe that all full-fidelity methods surpass the RS baseline (0.214) with the following errors: Vanilla BO (0.211), MIPROv2 (0.198), EASE (0.195), BOPCA (0.192), and HDBO (0.185). However, they all have higher error values than HbBoPs (0.150). Additionally, HbBoPs also outperforms all multi-fidelity methods. Although both TRIPLE-GSE (0.158) and TRIPLE-SH (0.159) exhibit superior performance compared to their best-in-class full-fidelity counterpart, i.e., HDBO (0.185), they on average have identified prompts that yield error values higher than the ones obtained for HbBoPs’s prompts.
Looking at the anytime performance with a more limited budget, e.g., a fraction of 0.25 LLM calls, we can confirm HbBoPs’s improvements over the baselines. Indeed, HbBoPs on average outperforms HDBO, the best full-fidelity method, by approximately 35% and TRIPLE-SH, the best multi-fidelity method, by 24%. For additional statistical analyses, we refer to Appendix E.
5.2 Analysis of the Performance for each LLM
Table 2: Median relative validation and test improvement of HbBoPs over TRIPLE-SH across ten benchmarks per LLM at different fractions of total LLM calls. IQR in parentheses.